OpenStudy (anonymous):

How would I expand the expression as a sum, difference, and/or constant multiple..... log7(xy^4/z^3) . I came up with 1/2[7logXY^4-7logz^3] not sure if I am on right track?

OpenStudy (campbell_st):

ok... here are a few laws you need to know about 1. \[\log(ab) = \log(a) + \log(b)\] 2. \[\log(x^a) = alog(x)\] and lastly \[\log(\frac{a}{b}) = \log(a) - \log(b)\] ok so you question is \[\log7(\frac{xy^4}{z^3})\] start by separating the 7 and the brackets... using the 1st log law shown above

OpenStudy (anonymous):

I see, so it would start with log7(xy^4)+log7(z^3) .. so far does that sound right?

OpenStudy (campbell_st):

not quite... I just want to check that its not a base 7 log is it... ? \[\log_{7}(\frac{xy^4}{x^3})\] is it..?

OpenStudy (anonymous):

ummm, the way you have problem typed is correct regarding the position of the 7 next to the log

OpenStudy (campbell_st):

ok.... choice 1 \[\log_{7}(\frac{xy^4}{z^3})\] or choice 2 \[\log7(\frac{xy^4}{z^3})\] which choice... is the question..?

OpenStudy (anonymous):

choice 1 ... sorry about not placing the numbers correctly, I am still new to the site and trying to figure out how to format questions correctly :)

OpenStudy (campbell_st):

ok... that makes it easier... I use the division rule 1st \[\log_{7} (xy^4) - \log_{7}(z^3)\] separate the term in the 1st log using the multiplication rule... rule 1?

OpenStudy (anonymous):

gotcha, so I am thinking \[\log_{7}(x)+\log_{7}(y ^{4})+\log_{7}(z ^{3}) \]

OpenStudy (campbell_st):

nearly... is \[\log_{7}(x) + \log_{7}(y^4) - \log_{7}(z^3)\] you just need to check your signs... last step... use the power rule... rule 2.. on the 2nd and 3rd logs

OpenStudy (anonymous):

ah I got the 3rd log mixed up but didn't need to do anything for the 2nd step. So the last step I am coming up with ... \[\log_{7}+4\log_{7}y-3\log_{7}z \]

OpenStudy (anonymous):

not sure if I need the () around the letters y&Z still?

OpenStudy (campbell_st):

nearly just fix the 1st log... you left out the x..

OpenStudy (campbell_st):

I normally write the term in brackets... it avoids confusion

OpenStudy (anonymous):

oh duh, lol Good point on the brackets so the 1st log should be \[\log_{7}(x) \] and of course the other logs am I correct?

OpenStudy (campbell_st):

yep \[\log_{7}(x) + 4\log_{7}(y) - 3\log_{7}(z)\]

OpenStudy (campbell_st):

good work.

OpenStudy (anonymous):

makes so much more sense the way you explained it :) I appreciate your time and explanation. Thank you!